The perimeter of an isosceles triangle is 37 in. The lengths of the two equal legs are 6 in. less than 3 times the length of the base. Find the lengths of the three sides.
Answer by ptaylor(2198)


Let x=length of base
And 3x=three times the length of the base
Then each of the two equal sides=3x-6
Now we are told that the perimeter (distance around the triangle)=37 in, so our equation to solve is:
x+(3x-6)+(3x-6)=37 get rid of parens
x+3x-6+3x-6=37 simplify
7x-12=37 add 12 to both sides
7x-12+12=37+12 collect like terms
7x=49 divide both sides by 7
x=7 in --------------------------------length of base
3x-6=3*7-6=15 in ---------------------length of each of the two equal sides
CK
15+15+7=37
37=37
Hope this helps----ptaylor